A mechanic has 20 liters of a 50% antifreeze solution. how many liters of pure antifreeze should be added to the antifreeze solution to make an antifreeze solution that is 75% antifreeze?
Answer by checkley77(12844)

.5*20+X=.75(20+X)
10+X=15+.75X
X-.75X=15-10
.25X=5
X=5/.25
X=20 LITERS OF PURE ANTIFREEZE IS USED.
PROOF:
.5*20+20=.75(20+20)
10+20=.75*40
30=30